Seven metrics

The following data were taken from the financial statements of Woodwork Enterprises Inc. for the current fiscal year. Assuming that there are no intangible assets.

Property, plant, and equipment (net) $ 5,000,000
Liabilities:
Current liabilities $ 400,000
Mortgage note payable, 5%, ten-year note issued two years ago 3,600,000
Total liabilities $4,000,000
Stockholders' equity:
Preferred $1 stock, $10 par (no change during year) $1,000,000
Common stock, $5 par (no change during year) 2,000,000
Retained earnings:
Balance, beginning of year $8,000,000
Net income 500,000 $8,500,000
Preferred dividends $ 100,000
Common dividends 100,000 (200,000)
Balance, end of year 8,300,000
Total stockholders' equity $11,300,000
Sales $ 6,250,000
Interest expense $ 180,000
Beginning-of-the-year amounts:
Property, plant, and equipment (net) $ 4,500,000
Total assets 12,200,000
Retained earnings 8,000,000

Determine the following: (a) debt ratio, (b) ratio of fixed assets to long-term liabilities, (c) ratio of liabilities to stockholders’ equity, (d) asset turnover, (e) return on total assets, (f) return on stockholders’ equity, and (g) return on common stockholders' equity. Round to two decimal places.

d. Asset turnover
e. Return on total assets %

Feedback:

(d) Divide net sales by average total assets, excluding long-term investments. Average Total Assets = (Beginning Total Assets + Ending Total Assets) ÷ 2
To find ending total assets, use the accounting equation and substitute ending liabilities + stockholders’ equity for the amount.

(e) Divide the sum of net income plus interest expense by average total assets. Average Total Assets = (Beginning Total Assets + Ending Total Assets) ÷ 2
To find ending total assets, use the accounting equation and substitute ending liabilities + stockholders’ equity for the amount.
Answer #1

Answer

--Requirement asked: 'D' and 'E'

--Working #1

A Total Stockholder's Equity $11,300,000
B Total Liabilities $4,000,000
C = A+B TOTAL ASSETS at the end $15,300,000

--Working #2

A Beginning Total Assets $12,200,000
B Ending Total Assets $15,300,000
C = (A+B)/2 Average Total Assets $13,750,000

--Requirement 'D"

A Sales [net Sales] $6,250,000
B [calculated above] Average Total Assets $13,750,000
C = A 'divided by' B Asset Turnover 0.454545455 or 0.45

--Requirement 'E'

A Net Income $500,000
B Interest Expense $180,000
C = A+B Net Income before Interest Expense $680,000
D [calculated above] Average Total Assets $13,750,000
E = C 'divided by' D Return On total assets 4.95%